\begin{thebibliography}{12}
\newcommand{\enquote}[1]{``#1''}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[\protect\citeauthoryear{Aliprantis and Border}{Aliprantis and
  Border}{2006}]{aliprantis2006infinite}
\textsc{Aliprantis, C.~D. and K.~Border} (2006): \emph{Infinite Dimensional
  Analysis: a hitchhiker's guide}, Springer.

\bibitem[\protect\citeauthoryear{Bertsekas}{Bertsekas}{2012}]{bertsekas2012dynamic}
\textsc{Bertsekas, D.} (2012): \emph{Dynamic programming and optimal control:
  Volume I}, vol.~1, Athena scientific.

\bibitem[\protect\citeauthoryear{Bertsekas and Shreve}{Bertsekas and
  Shreve}{1996}]{bertsekas1996stochastic}
\textsc{Bertsekas, D. and S.~E. Shreve} (1996): \emph{Stochastic optimal
  control: the discrete-time case}, vol.~5, Athena Scientific.

\bibitem[\protect\citeauthoryear{Blackwell}{Blackwell}{1965}]{blackwell1965discounted}
\textsc{Blackwell, D.} (1965): \enquote{Discounted dynamic programming,}
  \emph{The Annals of Mathematical Statistics}, 36, 226--235.

\bibitem[\protect\citeauthoryear{Dubins and Savage}{Dubins and
  Savage}{1965}]{dubins1965gamble}
\textsc{Dubins, L. and L.~Savage} (1965): \enquote{How To Gamble If You Must,}
  \emph{Mcgraw-Hill}.

\bibitem[\protect\citeauthoryear{Feinberg, Kasyanov, and Zadoianchuk}{Feinberg
  et~al.}{2012}]{feinberg2012average}
\textsc{Feinberg, E.~A., P.~O. Kasyanov, and N.~V. Zadoianchuk} (2012):
  \enquote{Average cost Markov decision processes with weakly continuous
  transition probabilities,} \emph{Mathematics of Operations Research}, 37,
  591--607.

\bibitem[\protect\citeauthoryear{Feinberg, Kasyanov, and Zgurovsky}{Feinberg
  et~al.}{2016}]{feinberg2016partially}
\textsc{Feinberg, E.~A., P.~O. Kasyanov, and M.~Z. Zgurovsky} (2016):
  \enquote{Partially observable total-cost Markov decision processes with
  weakly continuous transition probabilities,} \emph{Mathematics of Operations
  Research}, 41, 656--681.

\bibitem[\protect\citeauthoryear{Light}{Light}{2021}]{light2021stochastic}
\textsc{Light, B.} (2021): \enquote{Stochastic Comparative Statics in Markov
  Decision Processes,} \emph{Mathematics of Operations Research}, 46, 797--810.

\bibitem[\protect\citeauthoryear{Light}{Light}{2024}]{light2024course}
---\hspace{-.1pt}---\hspace{-.1pt}--- (2024): \enquote{A Course in Dynamic
  Optimization,} \emph{arXiv preprint arXiv:2408.03034}.

\bibitem[\protect\citeauthoryear{Maitra}{Maitra}{1968}]{maitra1968discounted}
\textsc{Maitra, A.} (1968): \enquote{Discounted dynamic programming on compact
  metric spaces,} \emph{Sankhy{\=a}: The Indian Journal of Statistics, Series
  A}, 211--216.

\bibitem[\protect\citeauthoryear{Puterman}{Puterman}{2014}]{puterman2014markov}
\textsc{Puterman, M.~L.} (2014): \emph{Markov decision processes: discrete
  stochastic dynamic programming}, John Wiley \& Sons.

\bibitem[\protect\citeauthoryear{Topkis}{Topkis}{1998}]{topkis1998supermodularity}
\textsc{Topkis, D.~M.} (1998): \emph{Supermodularity and complementarity},
  Princeton university press.

\end{thebibliography}

\documentclass[letter, 11pt]{article}

\usepackage[numbers]{natbib}
\usepackage{amssymb,amsmath,xcolor,graphicx,xspace,colortbl,rotating} %
\usepackage{subfigure}
\usepackage{appendix}  
\usepackage{bm} 
\usepackage{cancel} 
\usepackage{boxedminipage}  
\usepackage{color}  
\usepackage{endnotes}  
\usepackage{ragged2e}  
\usepackage[onehalfspacing]{setspace}  
\usepackage{tabulary}  
\usepackage{textcomp}  
\usepackage{varioref}  
\usepackage{graphicx}
\graphicspath{ {./images/} }
\usepackage[normalem]{ulem}
 
\usepackage[margin=1in]{geometry}
\newtheorem {theorem}{Theorem}

\newtheorem {lemma}{Lemma}

\newenvironment {proof}[1][Proof]{\noindent \textbf {#1.} }{\ \rule {0.5em}{0.5em}}

\makeatletter
\let\@fnsymbol\@arabic
\makeatother
\begin{document}

\newcommand{\newbl}[1]{{\color{blue}  #1}}
\newcommand{\ap}[1]{{\color{red} [Andres:  #1]}}

\title{The Principle of Optimality in Dynamic Programming: A Pedagogical Note}

\author{Bar Light\protect\thanks{Business School and Institute of Operations Research and Analytics, National University of Singapore, Singapore. e-mail: \textsf{barlight@nus.edu.sg} }}  
\maketitle

\thispagestyle{empty}

 \noindent \noindent \textsc{Abstract}:
\begin{quote}

The principle of optimality is a fundamental aspect of dynamic programming, which states that the optimal solution to a dynamic optimization problem can be found by combining the optimal solutions to its sub-problems. While this principle is generally applicable, it is often only taught for problems with finite or countable state spaces in order to sidestep measure-theoretic complexities. Therefore, it cannot be applied to classic models such as inventory management and dynamic pricing models that have continuous state spaces, and students may not be aware of the possible challenges involved in studying dynamic programming models with general state spaces. To address this, we provide conditions and a self-contained simple proof that establish when the principle of optimality for discounted dynamic programming is valid. These conditions shed light on the difficulties that may arise in the general state space case.  We provide examples from the literature that include the relatively involved case of universally measurable dynamic programming and the simple case of finite dynamic programming  where our main result can be applied to show that the principle of optimality holds. 

   
\end{quote}



\newpage

\section{Introduction} 

Dynamic programming is a powerful tool used in many areas, including operations research, engineering, artificial intelligence, mathematical finance,  and economics. One of the major concepts in dynamic programming is the principle of optimality (sometimes called Bellman's principle of optimality), which is a method for solving a dynamic optimization problem by  breaking it down into smaller sub-problems. While the principle of optimality holds under general conditions, proving it in the general case is quite involved and requires a non-trivial background in measure theory \citep{bertsekas1996stochastic}. In addition, some important functions that are studied in dynamic programming are not well behaved with respect to measurability. For example,   
it is well known, that even under the natural assumption that the dynamic optimization problem's primitives are Borel measurable, the value function need not be Borel measurable \citep{blackwell1965discounted}. 

To overcome the measurability difficulty, major textbooks that are used to teach the principle of optimality in dynamic programming, even for advanced graduate level courses, typically avoid measure-theoretic issues by assuming that the state space is finite or countable (e.g., \cite{bertsekas2012dynamic} and \cite{puterman2014markov}). This approach has two major drawbacks. First, in some of the canonical dynamic optimization models in operations research and economics like inventory management, consumption-savings problems, or dynamic pricing models, the state space is  continuous. Hence, the theory that is developed for countable state spaces does not apply to these canonical models. 
Second, students who are introduced to the important field of dynamic programming may not fully comprehend or acknowledge the challenges involved in studying dynamic programming models with continuous or other more  general state spaces such as Borel spaces.

In this note, we provide a self-contained proof for the principle of optimality and the existence of an optimal stationary policy for discounted  dynamic programming that requires minimal knowledge of measure theory. The key assumptions that guarantee the correctness of the principle of optimality and the existence of an  $\epsilon$-optimal stationary policy are that the Bellman operator preserves  
 some form of measurability and that there is a measurable selection that attains the supremum of the Bellman operator. These assumptions are simple to state and provide insight into the potential measurability challenges in dynamic programming with general state spaces (such as the issue of measurable selection).  
 What differentiates the paper's approach from previous literature is the nature of these assumptions. Unlike typical results, where assumptions are made directly on the primitives of the dynamic programming model (e.g., continuity or specific forms of measurability), the paper's assumptions are instead placed on the properties of the Bellman operator itself. This allows us to provide a relatively simple proof that is similar to the proof for the finite state space case, but our result can be applied to much more involved cases, as demonstrated through  our examples. Importantly, while our assumptions hold trivially for finite or countable state spaces, establishing them in the general case can be quite involved. Typically, proving these assumptions relies on pure measure-theoretic or topological arguments that are somewhat disconnected from the essence of dynamic programming itself. Hence, we provide a framework that is both accessible and helps to bridge the gap between simple and involved cases without straying  from the core concepts of dynamic programming.

  Our results are applicable to important cases that have been studied in the literature.  We provide seven examples from the literature where our main result can be applied to prove the principle of optimality and the existence of an optimal policy, including the general universally measurable case \citep{bertsekas1996stochastic}, the upper semi continuous case 
 \citep{maitra1968discounted}, the monotone case \cite{topkis1998supermodularity,light2021stochastic}, and the standard countable case  \citep{bertsekas2012dynamic}. We demonstrate a potential failure of our assumptions using an example from \cite{blackwell1965discounted} that involves Borel measurable primitives. The rest of the note is organized as follows.  In Section 2 we present the discounted dynamic programming model. In Section 2.1 we present the main theorem and its proof. In Section 2.2 we present the applications of the main Theorem.

\section{Discounted Dynamic Programming}

For the sake of brevity, we will focus on discounted dynamic programs in this note.

We define a discounted dynamic programming model in terms of a tuple of elements $(S ,A ,\Gamma  ,p , U , r , \beta  )$. 
The state space $S$ is a Borel measurable separable metric space that describes the possible states of the system. The action space $A$ is a Borel measurable separable metric space. We denote by $\mathcal{B}(X)$ the Borel sigma-algebra on a Polish space $X$.

The set $U:=U(X,Y)$ is a set of functions from a Polish space $X$ to a Polish space $Y$, $f:X \rightarrow Y$,  such that $f$ is $\mathcal{U}$-measurable in the sense that $f^{-1}(B) \in \mathcal{U}(X)$ for all $B \in \mathcal{B}(Y)$,  where $\mathcal{U}(X)$ is a sigma-algebra that is well defined for every Polish space. We will assume in this note that $\mathcal{U}(X)$ consists of the Borel  measurable sets, the universally measurable sets, or all sets on $X$. We choose to focus on these sets because they are widely studied in the literature on dynamic optimization (see Section \ref{Sec:applications}). We can study other measurable sets in our framework as long as certain properties are satisfied to ensure that the Bellman operator preserves measurability (see Lemma \ref{lemma:T-lambda-closed} for more details).    Thus, $U(X,Y)$ can be the set of all Borel measurable functions, universally measurable functions, or all functions from $X$ to $Y$. The function 
$p :S \times A \times \mathcal{B}(S) \rightarrow [0 ,1]$ is a transition probability function, i.e., $p(s ,a , \cdot)$ is a probability measure on $(S,\mathcal{B}(S))$ for each $(s ,a) \in S \times A$ and $p(\cdot,\cdot,B)$ is a $\mathcal{U}$-measurable function for each $B \in \mathcal{B}(S)$. The function 
$r : S \times A \rightarrow \mathbb{R}$ is the single-period 
  payoff function that is bounded from below, and $\mathcal{U}$-measurable. $0 <\beta  < 1$ is the discount factor. The constraint set  $\Gamma$ is Borel measurable subset of $S \times A$ such that for any $s \in S$ the $s$-section $\Gamma(s)$ of $\Gamma$ is a  non-empty Borel measurable set of feasible actions in state $s \in S$.

The process starts at some state $s(1) \in S$. Suppose that at time $t$ the state is $s(t)$. Based on $s(t)$, the decision maker (DM) chooses an action $a(t) \in \Gamma(s(t))$ and receives a payoff $r(s(t) ,a(t))$. The probability that the next period's state $s(t +1)$ will lie in $B \in \mathcal{B}(S)$ is given by $p(s(t) ,a(t) ,B)$.

 Let $H^{t}:= S \times (A \times S)^{t-1} $ be the set of finite histories up to time $t$. A strategy $\sigma$ is a sequence $(\sigma _{1} ,\sigma _{2} , \ldots )$ of  functions $\sigma _{t} \in U(H^{t},A)$ such that $\sigma_{t}(s(1),a(1),\ldots,s(t)) \in \Gamma(s(t))$ for all $t$ and all $(s(1),a(1),\ldots,s(t)) \in H^{t}$. Each initial state $s(1) \in S$, a strategy $\sigma $, and a transition probability function $p$ induce a well-defined probability measure over the measurable space of all infinite histories $(H^{\infty },\mathcal{B}(H^{\infty }))$ (see Propositions 7.28 and 7.45 in \cite{bertsekas1996stochastic}). 
 We denote the expectation with respect to that probability measure by $\mathbb{E}_{\sigma }$, and the associated stochastic process by $\{s(t) ,a(t)\}_{t =1}^{\infty }$. 
The decision maker's goal is to find a strategy that maximizes her expected discounted payoff. When the decision maker follows a strategy $\sigma$ and the initial state is $s \in S$, the expected discounted payoff is given by
\begin{equation} \label{eq: V_sigma}
V_{\sigma}(s) =\mathbb{E}_{\sigma }\sum \limits_{t =1}^{\infty }\beta^{t -1}r(s(t),a(t)).
\end{equation}
Define 
\begin{equation*} 
V(s) =\sup _{\sigma }V_{\sigma }(s) .
\end{equation*}
We call $V :S \rightarrow \mathbb{R}$ the value function. A strategy $\sigma$ is said to be optimal if $V_{\sigma} (s) = V(s)$ for all $s$. A strategy $\sigma$ is said to be $\epsilon$-optimal if for every $\epsilon >0$, every strategy $\sigma'$, and any $s\in S$ we have $V_{\sigma'}(s) \leq V_{\sigma}(s) + \epsilon$. A strategy is said to be stationary if $\sigma (h(t)) = \lambda(s(t))$ where $h(t) = (s(1),a(1),\ldots,s(t) )$ for any history $h(t) \in H^{t}$ and some function $\lambda \in U(S,A)$ which is called a stationary policy. 

For a function $f \in U(S,\mathbb{R})$ we define the operator (called the Bellman operator) $T$ by 
\begin{equation*}Tf(s) =\sup _{a \in \Gamma(s)}Q(s ,a ,f), 
\end{equation*}
where 
\begin{equation}Q(s ,a ,f) = r(s ,a) +\beta \int _{S}f(s^{ \prime })p(s ,a ,ds^{ \prime }) \label{eq:h}
\end{equation}
and  the set of optimal solutions for the optimization problem above $\Lambda _{f}$ by
$$\Lambda _{f} = \{ \lambda \in U(S,A): \sup _{a \in \Gamma(s)} \ Q(s,a,f) = Q(s,\lambda(s),f), \  \forall s \in S \} .$$

 For a stationary policy $\lambda \in U(S,A)$ and $f \in U(S,\mathbb{R})$ we also define $T_{\lambda}$ by
\begin{equation}
    T_{\lambda}f(s) =  r(s,\lambda(s)) + \beta \int _{S} f(s') p(s,\lambda(s),ds') = Q(s,\lambda(s),f).
\end{equation}

\subsection{The Dynamic Programming Principle} 

Let $B(S)$ be the space  of all real-valued bounded  functions on $S$ equipped with the sup-norm $ \Vert f \Vert = \sup _{s \in S} |f(s)| $ and the induced metric $d_{\infty} (f,g):= \sup _{s \in S} |f(s) - g(s)|$  for $f,g \in B(S)$. Let $B_{m}(S) =  B(S) \cap U(S,\mathbb{R})$.  We now present the main theorem of this note that shows that the value function satisfies the Bellman equation under certain conditions.  We show that if there exists a closed set $D$ of $B_{m}(S)$ such that $T(D) \subseteq D$ and for each $f \in D$ we have $\Lambda_{f} \neq \emptyset$, then the value function is in $D$, satisfies the Bellman equation, and there exists an $\epsilon$-optimal stationary policy. We note that we could also prove a version of Theorem \ref{Thm:mainDP} for the case of unbounded but positive rewards (discounted or undiscounted) by using a monotone convergence theorem argument instead of Banach's fixed-point theorem (see \cite{bertsekas1996stochastic}, \cite{feinberg2012average} and \cite{feinberg2016partially}).

The assumption that $f \in D$ implies that $Tf \in D$  guarantees that the Bellman equation preserves some form of measurability as $D$ consists of appropriately chosen measurable functions. The assumption that $f \in D$ implies that  $\Lambda_{f} \neq \emptyset$ guarantees a measurable solution (or a measurable selection) for the Bellman equation. In the finite and countable cases these assumptions trivially hold (see Section \ref{Sec:applications}). In  the general case, these assumptions can be quite involved to establish. Finding  
 conditions on the model's primitives that guarantee that these assumptions hold are crucial in the study of dynamic programming. We discuss some important cases from the literature in Section \ref{Sec:applications}.

\begin{theorem} \label{Thm:mainDP} Let $D \subseteq B_{m}(S)$ be a complete metric space with the metric $d_{\infty}$ that contains the constant $0$ function and any constant function $M$ such that $M= \sup_{s \in S} f(s)$ for $f \in D$. 

Assume that $f \in D$ implies $\Lambda_{f} \neq \emptyset$, and $Tf \in D$.  

Then the following holds: 

1. $V$ is the unique function in $D$ that satisfies the Bellman equation, i.e., $TV = V$.

2. There is a stationary policy $\lambda \in \Lambda_{V}$ that is $\epsilon$-optimal. 

3. If $\lambda \in \Lambda_{V}$ achieves the supremum then $\lambda$ is an optimal stationary policy. 

\end{theorem}

Note that  $U(S,\mathbb{R})$ is closed in $B(S)$ so $B_{m}(S)$ is a complete metric space.

We first prove a few simple properties of the operators $T$ and $T_{\lambda}$. 

\begin{lemma} \label{lemma:T-lambda-closed}
     For each $\lambda \in U(S,A)$, $f \in U(S,\mathbb{R})$ implies $T_{\lambda}f \in U(S,\mathbb{R})$. In addition, for each strategy $\sigma$, we have $V_{\sigma} \in U(S,\mathbb{R})$.   
\end{lemma}
\begin{proof}
Let $f \in U(S,\mathbb{R})$ and $\lambda \in U(S,A)$. 
The fact that $T_{\lambda} f \in U(S,\mathbb{R})$ follows immediately for the case that $U$ is the set of all functions. For the other cases, it holds from the following argument:
 $\int _{S} f(y) p(s,a,ds')$ is universally measurable (Borel measurable) whenever $f$ is universally measurable (Borel measurable). Hence,  $Q(s,a,f) $ is universally measurable  (Borel measurable) as the sum of universally measurable (Borel measurable) functions (recall that $r$ is assumed to be $\mathcal{U}$-measurable). Thus, $T_{\lambda}f = Q(s,\lambda(s),f)  $ is universally measurable (Borel measurable) as the composition of  universally measurable (Borel measurable) functions (see Chapter 7 in \cite{bertsekas1996stochastic} for a proof of these properties of Borel and universally measurable functions). 

Using the same argument as above repeatedly, and because, $U$ is closed under composition, summation, and pointwise limits, $V_{\sigma}$ is $\mathcal{U}$-measurable for each strategy  $\sigma$. 
\end{proof}

For two functions $f,g$ from $S$ to $\mathbb{R}$, $f \geq g$ means $f(s) \geq g(s)$ for each $s \in S$.
\begin{lemma} \label{lemm:Tf-prop}
The operator $T$ satisfies the following two properties.

(a) Monotone: $Tf \geq Tg$ whenever $f \geq g$. 

(b) Discounting: $T(f+c) = Tf + \beta c$ for any constant function $c$. 

In addition, $T_{\lambda}$ satisfies $(a)$ and $(b)$ for any stationary policy $\lambda \in U(S,A)$. 
\end{lemma}

\begin{proof}
    (a) We have $Q(s,a,f) \geq Q(s,a,g)$ whenever $f \geq g$. Hence, $Tf \geq Tg$.

    (b) For each $s \in S$ we have
    $$T(f+c)(s) =  \sup _{a \in \Gamma(s)} r(s ,a) +\beta \int _{S}f(s^{ \prime })p(s ,a ,ds^{ \prime }) + \beta c = Tf(s) + \beta c.$$

   The proof for the operator $T_{\lambda}$ is similar. 
\end{proof}

A key property of the operators $T$ and $T_{\lambda}$ is the contraction property. The operator $T$ on $B_{m}(S)$ is $L$-contraction if $\Vert Tf - Tg \Vert \leq L \Vert f - g \Vert$ for all $f,g \in B_{m}(S)$. 

\begin{lemma} \label{lemm:Tf-contraction}
    The operator $T$ is $\beta$-contraction on $B_{m}(S)$, and the operator $T_{\lambda}$ is $\beta$-contraction on $B_{m}(S)$ for any stationary policy $\lambda \in U(S,A)$. 
\end{lemma}

\begin{proof}
Let $f,g \in B_{m}(S) = U(S,\mathbb{R}) \cap B(S)$ and  $\lambda \in U(S,A)$. From  Lemma \ref{lemma:T-lambda-closed}, $T_{\lambda}f \in U(S,\mathbb{R})$ and from the argument in the start of the proof of Theorem \ref{Thm:mainDP}, $T_{\lambda}f \in B(S)$ so $T_{\lambda}f$ is indeed in $B_{m}(S)$. 

For any $s \in S$ we have 
$$ | T_{\lambda}f(s) - T_{\lambda} g(s) | = |\beta \int_{S}(f(s') - g(s'))p(s,\lambda (s),ds') | \leq \beta \int _{S} | f(s') - g(s')|p(s,\lambda(s),ds') \leq \beta \Vert f - g \Vert. $$
Taking the supremum over $s \in S$ yields $\Vert T_{\lambda}f - T_{\lambda} g \Vert \leq \beta \Vert f - g \Vert$ so $T_{\lambda}$ is  $\beta$-contraction on $B_{m}(s)$.

Hence, it follows that for all $s \in S$ and $a \in \Gamma(s)$ we have 
$$Q(s,a,f)  \leq Q(s,a,g) + \beta \Vert f - g \Vert  \leq Tg (s)+ \beta \Vert f - g \Vert.  $$
Thus, taking the supremum over $a \in \Gamma(s)$ yields 
$$ Tf(s) \leq Tg(s) + \beta \Vert f - g \Vert.$$
By a symmetrical argument, we can deduce $Tg(s) \leq Tf(s) + \beta \Vert f - g \Vert$ for all $s \in S$ so $T$ is $\beta$-contraction on $B_{m}(S)$. 
\end{proof}

\begin{proof} [Proof of Theorem \ref{Thm:mainDP}]
From the Banach fixed-point theorem (see Theorem 3.48 in \cite{aliprantis2006infinite}) and Lemma \ref{lemm:Tf-contraction} the operator $T:D \rightarrow D$ has a unique fixed-point $x$. We will show that $V=x$.  We first show that $V \leq x$. 

Let $ k (s) = \sup_{a \in \Gamma(s)} r(s,a)  $. Then $k(s)$ is finite. This follows because $Tf \in B_{m}(S)$ whenever $f$ is in $D$, so taking $f$ to be the $0$ constant function that belongs to $D$ by assumption implies that $M' \geq Tf(s) \geq r(s,a)$ for all $(s,a)$ for some finite $M'$. In addition, $Tf(s) = \sup _{a \in \Gamma(s)} r(s,a) =k (s)$ is in $D$. Thus, $ M = \sup_{s \in S} k(s) $ belongs to $D$ also.  Using the same argument again for the constant function $f_{2}(s) = M$ for all $s \in S$ we have $Tf_{2}(s) = k(s) + \beta M $ is in $D$ so $M + \beta M$ is in $D$. Continuing inductively, we conclude that the constant function $M_{T}$ given by $ M_{T} =\sum _{t=1}^{T} \beta ^{t-1} M$ is in $D$. Because $D$ is closed as a closed subset of a complete metric space, the constant function $ y(s)= \lim_{T \rightarrow \infty} M_{T} = M /(1-\beta)$  for all $s \in S$ belongs to $D$.  Because $r$ is bounded by $M$ we have $ V \leq y$. Hence, there exists a $y \in D $ such that $V \leq y$.

 For a strategy $\sigma$ let $\sigma | (s,a)$ be the strategy that is induced by $\sigma$ given that the first period's state-action pair was $(s,a)$, i.e., $\sigma | (s,a) _{t} (h_{t}) = \sigma _{t+1} (s,a,h_{t}) $ for all $t$ and any history $h_{t} \in H^{t}$.  
For any strategy $\sigma$, a function $y \in D$ with $V \leq y$, and $s \in S$ we have 
\begin{align*}
    V_{\sigma}(s) & = \mathbb{E}_{\sigma } \left ( r(s,a(1)) + \sum \limits_{t =2}^{\infty }\beta^{t -1}r(s(t),a(t)) \right )   \\
    & = r(s,\sigma_{1}(s)) + \beta \int _{S} V_{\sigma | (s,\sigma_{1}(s))} (s') p(s, \sigma_{1}(s),ds') \\
    & \leq  r(s,\sigma_{1}(s)) + \beta \int _{S} y (s') p(s, \sigma_{1}(s),ds') \\
    & \leq \sup _{a \in \Gamma(s)} \ r(s,a) + \beta \int _{S} y (s') p(s, a,ds') \\
    & = Ty(s). 
\end{align*}
   Thus, taking the supremum over all strategies $\sigma$ implies that $V \leq Ty$.  Repeating the same argument, we deduce that $V \leq T^{n}y$ for all $n \geq 1$. From the Banach fixed-point theorem $T^{n}y$ converges to the unique fixed-point of $T$ on $D$. We conclude that $V \leq x$.  

We now show that $V \geq x$. 
   Let $\epsilon >0$. Because $x \in D$, the Theorem's assumption implies that $\Lambda_{x} \neq \emptyset $. Hence, because $Tx=x$, there is a function $\lambda \in U(S,A)$ such that $T_{\lambda}x \geq x - \epsilon(1-\beta) $. 
   
   We will now show that $T_{\lambda}^{n}x \geq x - \epsilon$ for all $n$. For $n=1$ the result holds from the construction of $\lambda$. Assume it holds for $n>1$. Then 
   $$ T_{\lambda}^{n+1}x = T_{\lambda} (T_{\lambda}^{n}x) \geq T_{\lambda} (x - \epsilon) = T_{\lambda}x - \beta \epsilon \geq  x - \epsilon$$
   where we used Lemma \ref{lemm:Tf-prop} to derive the inequalities. 
Thus, $T_{\lambda}^{n}x \geq x - \epsilon$ for all $n$. From Lemma \ref{lemm:Tf-contraction}  and the Banach fixed-point theorem  
 $T_{\lambda}^{n}$ converges to the unique fixed-point of $T_{\lambda}$ in $B_{m}(S)$ which we denote by $W$ so $W \geq x - \epsilon$. So $W:S \rightarrow \mathbb{R}$ is $\mathcal{U}$-measurable and we have 
$$W (s) = r(s,\lambda(s)) + \beta \int_{S}W(s')p(s,\lambda(s),ds') = \mathbb{E}_{\sigma }\sum \limits_{t =1}^{\infty }\beta^{t -1}r(s(t),a(t)) $$
with $s=s(1)$ and $\sigma$ is the stationary plan that plays according to $\lambda$, i.e., $\sigma_{t}(h_{n}) = \lambda(s(n))$ for all $h_{n}(s(1),a(1),\ldots,s(n))$. Hence, $V \geq W \geq x - \epsilon$. Thus, $V \geq x$.  We conclude that $V=x$ so $V$ is the unique fixed point of $T$ on $D$ and $\lambda \in \Lambda_{V}$ is an $\epsilon$-optimal strategy.  

In addition, if $\lambda \in \Lambda_{V}$ achieves the supremum then 
$T_{\lambda} V = V$ so $\lambda$ is the optimal stationary policy.
\end{proof}

Under the conditions of Theorem \ref{Thm:mainDP}, $V$ belongs to the set $D$.  Hence, we can use Theorem \ref{Thm:mainDP} to prove properties of the value function such as monotonicity or concavity as we show in Section \ref{Sec:applications}.

\subsection{Applications of Theorem \ref{Thm:mainDP}}    \label{Sec:applications}
As a starting point, it is natural to assume that the dynamic programming problem's primitives  are Borel measurable and  choose $D$ and $U$ to be the sets of Borel measurable functions. But there is a well-known measurability issue with this choice. The following example with Borel primitives is provided by \cite{blackwell1965discounted}. 
Let $S=A=[0,1]$. Then there is a Borel measurable set $B \subseteq S \times A$ such that its projection $E$ to $S$ is not Borel measurable.  Consider $f$ to be the zero function and $r = 1_{B}(a,s)$. Then $f$ and $r$ are Borel measurable. But $Tf (s) = \sup_{a \in A} 1_{B}(a,s) = 1_{E}(s)$ is not Borel measurable. Hence, we cannot apply Theorem \ref{Thm:mainDP} for this case. 

We now present the assumptions for some of the main cases studied in the literature on discounted dynamic programming, where Theorem 
\ref{Thm:mainDP} can be applied.

We say that $p$ is continuous  (concave, increasing, Borel measurable) if $\int_{S} f(s')p(s,a,ds')$ is continuous (concave, increasing, Borel measurable) on $S \times A$ for every bounded and continuous (concave, increasing, Borel measurable) function $f$ on $S$. 
\begin{itemize}
    \item  \textbf{The upper semianalytic case.} The set  $U(X,Y)$ is the set of universally measurable functions from $X$ to $Y$ (see \cite{bertsekas1996stochastic} for a detailed study and definitions of analytic sets, upper semianalytic functions, and universally measurable sets and functions).   The payoff function $r$ is a bounded upper semianalytic function, and $p$ is Borel measurable. We choose $D$ to be the set of all bounded upper semianalytic functions on $S$. The set $D$ is closed in $B_{m}(S)$ and it can be shown that $f \in D$ implies $Tf \in D$ and $\Lambda_{f} \neq 0$ (see Propositions 7.47 and 7.5 in \cite{bertsekas1996stochastic}). That is, $Tf$ is upper semianalytic and there exists a universally measurable selection. Hence, we can apply Theorem \ref{Thm:mainDP}.  

For the next example, recall that a function $f:S \rightarrow \mathbb{R}$ is upper semi continuous (u.s.c) if $\limsup f(s_{n}) \leq f(s)$ whenever $s_{n} \rightarrow s$.  The correspondence $\Gamma:S \rightarrow 2^{A}$ is u.s.c  if $\{s \in S: \Gamma(s) \subseteq U \}$ is an open set whenever $U$ is an open set in $A$ and lower semi continuous (l.s.c) if $\{s \in S: \Gamma(s) \cap U \neq \emptyset \}$ is an open set whenever $U$ is an open set in $A$ (see Chapter 17 in \cite{aliprantis2006infinite}).

\item \textbf{The upper semi continuous case.} The set $U(X,Y)$ is the set of Borel measurable functions from $X$ to $Y$. The set $\Gamma(s)$ is a compact for each $s \in S$, $\Gamma$ is u.s.c, $r$ is bounded and u.s.c, and $p$ is continuous. Under these conditions, we can choose $D$ to be the set of all bounded u.s.c functions on $S$. The set $D$ is closed in $B_{m}(S)$ and it can be shown that $f \in D$ implies $Tf \in D$ and $\Lambda_{f} \neq 0$ (see \cite{dubins1965gamble} and \cite{maitra1968discounted} for proofs of these claims and Section 2.3 in \cite{light2024course} for a textbook level treatment of the upper semi continuous case). In particular, the value function $V$ is u.s.c and there is a Borel measurable selection that achieves the supremum. Hence, we can apply Theorem \ref{Thm:mainDP}.

\item \textbf{The continuous case.} The set  $U(X,Y)$ is the set of Borel measurable functions from $X$ to $Y$. The set $\Gamma(s)$ is a compact   for each $s \in S$, $\Gamma$ is u.s.c and l.h.s, $r$ is bounded and continuous, $p$ is continuous. We can choose $D$ to be the set of all bounded and continuous functions on $S$ which is closed in $B_{m}(S)$. It can be shown that if $f \in D$ then $Tf \in D$  (see Theorem 17.31 in \cite{aliprantis2006infinite}) and $\Lambda_{f} \neq \emptyset$ from the u.s.c case. In particular, the value function $V$ is continuous. Hence, we can apply Theorem \ref{Thm:mainDP}.

The correspondence $\Gamma$ is called convex if  $a_{1} \in \Gamma (s_{1})$ and $a_{2} \in \Gamma (s_{2})$ imply that $a_{\lambda} \in \Gamma (s_{\lambda})$ for all $s_{1},s_{2}$ and $\lambda \in (0,1)$ where   
 $s_{\lambda} = \lambda s_{1} +(1-\lambda) s_{2}$ and $a_{\lambda} = \lambda a_{1} +(1-\lambda) a_{2}$.
 
\item \textbf{The concave case.} The set $U(X,Y)$ is the set of Borel measurable from $X$ to $Y$. The set $S$ is convex, $\Gamma$ is convex, $r$ is concave and is strictly concave in $a$, $p$ is concave, and all the assumptions of the u.s.c case hold. We can choose $D$ to be the set of all bounded, u.s.c and concave functions on $S$ which is closed in $B_{m}(S)$.   It can be shown that if $f \in D$ then $Tf \in D$ (see the u.s.c case and Theorem 3.1 in \cite{light2024course} for concavity)  and $\Lambda_{f} \neq \emptyset$ from the u.s.c case. In particular, the value function $V$ is concave and the optimal policy is single-valued. Hence, we can apply Theorem \ref{Thm:mainDP}.  

We need the following definitions for the monotone case. A lattice $X$ is a partially ordered set such that each two elements $x,y \in X$ has a least upper bound (denoted by $x \lor y$) and a greatest lower bounded (denoted by $x \land y$). 
Recall that a function $f:X \rightarrow \mathbb{R}$, where $X$ is a lattice in $\mathbb{R}^{n}$, (we assume the standard product order on $\mathbb{R}^{n}$, i.e., $x \geq y$, $x,y \in \mathbb{R}^{n}$ if $x_{i} \geq y_{i}$ for each $i=1,\ldots,n$) is supermodular if $f (x \vee y) + f(x \wedge y) \geq f(x) + f(y) $ for all $x,y \in X$ where as usual $x \vee y$ denotes the componentwise maximum and $x \wedge y$ the componentwise minimum. A correspondence $\Lambda$ from $X$ to $E$ where $\Lambda$ is a (sub)-lattice for each $x$, is ascending if $\Lambda (x_{2}) \succeq_{set} \Lambda(x_{1})$ whenever $x_{2} \geq x_{1}$ where for two sets $D$ and $E$, write $D \succeq _{set} E$ if $x \in D$ and $x' \in E$ imply $x \lor x' \in D$ and $x \land x' \in E$ (see \cite{topkis1998supermodularity} for a comprehensive treatment of supermodularity). 
\item \textbf{The monotone case.} (i) (Monotone value function). The set $U(X,Y)$ is the set of Borel measurable functions  from $X$ to $Y$. The set $A \subseteq \mathbb{R}^{n}$ and $S\subseteq \mathbb{R}^{m}$. We have $\Gamma(s_{1}) \subseteq \Gamma(s_{2})$ whenever $s_{1} \leq s_{2}$, $r$ is increasing in $s$, $p$ is increasing in $s$, and all the assumptions of the u.s.c case hold. We can choose $D$ to be the set of all u.s.c, bounded, increasing functions which is closed in $B_{m}(S)$.  It can be shown that if $f \in D$ then $Tf \in D$ (see Theorem 3.92 in  \cite{topkis1998supermodularity} and Theorem 3.1 in \cite{light2024course})  and $\Lambda_{f}$ is non-empty from the u.s.c case. In particular, the value function is increasing. Hence, we can apply Theorem \ref{Thm:mainDP}.

(ii) (Monotone policy function). The set $U(X,Y)$ is the set of Borel measurable functions  from $X$ to $Y$. The sets $A \subseteq \mathbb{R}^{n}$ and $S\subseteq \mathbb{R}^{m}$ are lattices. The graph of $\Gamma$, $\operatorname{Gr} \Gamma = \{ (s,a) \in S \times A : a \in \Gamma(s) \}$ is a sub-lattice in $S \times A$, $r$ is supermodular, $p$ is supermodular, and all the assumptions of the u.s.c case hold. We can choose $D$ to be the set of all u.s.c, bounded, and supermodular functions which is closed in $B_{m}(S)$.  It can be shown that if $f \in D$ then $Tf \in D$, $\Lambda_{f}$ is non-empty, and the set of optimal solutions  $$G(s) = \operatorname{argmax} _{a \in \Gamma(s)} r(s,a) + \beta \int V(s')p(s,a,ds')$$ is ascending (see Theorem 3.4 in \cite{light2024course} and see Theorem 3.92 in  \cite{topkis1998supermodularity} and Theorem 3 in \cite{light2021stochastic} for related results).  In particular, the value function $V$ is supermodular. Hence, we can apply Theorem \ref{Thm:mainDP}.

\item \textbf{The countable-compact case}. The set $U(X,Y)$ is the set of all Borel measurable functions from $X$ to $Y$. The set $S$ is countable, $\Gamma(s)$ is compact in $\mathbb{R}^{n}$ (a prominent choice is $\Gamma(s) = A $ for each $s$,  and $A$ is the set all probability measures  on some finite action set $A'$), $r$ is  bounded. We can choose $D=B_{m}(S)=B(S)$ and the conditions of Theorem \ref{Thm:mainDP} obviously hold. 

\item \textbf{The finite case.} The set $U(X,Y)$ is the set of all functions from $X$ to $Y$. The sets $S$ and $A$ are finite.  We can choose $D=B_{m}(S)=B(S)$ and the conditions of Theorem \ref{Thm:mainDP} hold. 

\end{itemize}

\bibliographystyle{ecta}
\bibliography{DP}

\end{document}